OpenStudy (anonymous):

Factor the following polynomial completely 208p^2-325q^2

OpenStudy (cwrw238):

208 = 2 * 2 * 2 * 2 * 13 so sqrt208 = 4sqrt13 325 = 5 * 5 * 13 = 5sqrt13 so sqrt325 = 5sqrt13 208p^2-325q^2 = (4sqrt3p + 5sqrt13q)(4sqrt3p - 5sqrt13q)

OpenStudy (cwrw238):

i used the identity a^2 - b^2 = (a + b)(a - b)